OpenStudy (lyssburleson):

The sum of two numbers is 52 and the difference is 16. What are the numbers?

OpenStudy (imstuck):

You have to set this up in 2 different equations, using 2 different unknowns. The 2 numbers can be x and y. What does "sum" mean?

OpenStudy (mathmale):

Using "x" and "y," and the sum value 52, write an equation for "the sum of two numbers."

OpenStudy (alyssarayne03):

"The sum of two numbers is 52," x + y = 52 y = (52-x)

OpenStudy (imstuck):

good!

OpenStudy (lyssburleson):

so i just have to add something to 16 to get the sum which is 52?

OpenStudy (mathmale):

Not quite. You are to come up with 2 separate equations, and then to solve these equations simultaneously to obtain values for the variables x and y. x+y-52 Write an equation for "the difference between the two numbers is 16."

OpenStudy (mathmale):

sum: addition x+y=52 difference: subtraction: write the equation here. Solve this system of linear equations, preferably by addition / subtraction.